OpenStudy (anonymous):

What is the measure of side AC 8, 7, 9, 136, or none of the above

OpenStudy (anonymous):



jhonyy9 (jhonyy9):

so this a right triangle ,yes ?

OpenStudy (anonymous):

yep

jhonyy9 (jhonyy9):

then you need to know the pythagorean numbers so like in this case you can calculi it let the unknow being x so x^2 = 10^2 -6^2 can you calculi x ?

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

i'll see what I get

jhonyy9 (jhonyy9):

what will be x ?

OpenStudy (anonymous):

100=36-36

jhonyy9 (jhonyy9):

36-36 not is equal 100

OpenStudy (anonymous):

oops

jhonyy9 (jhonyy9):

100 need being equal 36 + x^2 so then x^2 = 100-36 = 64 then x^2 =64 so x = sqrt64 = ?

OpenStudy (anonymous):

8

jhonyy9 (jhonyy9):

exactly very nice

jhonyy9 (jhonyy9):

was easy yes ?

OpenStudy (anonymous):

yes, now the last thing you wrote really made sense.. thanks!

jhonyy9 (jhonyy9):

np was my pleasure

jhonyy9 (jhonyy9):

good luck bye

OpenStudy (anonymous):

you too... bye